# Bayesian Sheaf Neural Networks

### v1

@@ 5.1 Results
Though the results are mixed in the case of diagonal restriction maps, the $SO(d)$-BSNN outperforms the corresponding $O(d)$-NSD model on all three datasets, and the Gen-BSNN significantly outperforms the corresponding Gen-NSD on all three dataset. This suggests that the comparative advantage of our Bayesian approach to learning cellular sheaves is more pronounced for sheaves with a greater number of parameters. We also note that on each of the three WebKB datasets, the best test accuracy out of all the models is attained by one of our BSNNs. Moreover, the standard deviation of the reported test accuracies for each BSNN is lower than the corresponding NSD model with the same type of restriction map in all but one case, showing that our BSNN models are less sensitive to the effects of random weight initializations.
@@
From Table 2, the advantage of each BSNN model over the corresponding NSD model when considering the entire range of hyperparameters is definitive. For every type of restriction map and dataset, a Wilcoxon signed-rank test for a $p$-value threshold of $p=0.05$ indicates a statistically significant difference between the results of the BSNN model and the corresponding NSD model. Thus we observe that the BSNN models maintain good performance over a much wider range of hyperparameters when compared with the NSD model with the same type of restriction maps.
Table 1: Average test accuracies and standard deviations across 30 random seeds for the hyperparameter configuration yielding the highest validation accuracy for each model and dataset.
Texas Wisconsin Cornell
Diag-BSNN $75.35{\scriptstyle\pm 1.68}$ $\mathbf{83.04}{\scriptstyle\pm 0.87}$ $72.23{\scriptstyle\pm 1.76}$
Diag-NSD $\mathbf{75.49}{\scriptstyle\pm 2.40}$ $82.77{\scriptstyle\pm 1.15}$ $\textbf{{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}73.70}}{\scriptstyle\pm 1.45}$
$\mathbf{SO(d)}$-BSNN $\mathbf{76.65}{\scriptstyle\pm 1.48}$ $\mathbf{82.02}{\scriptstyle\pm{1.06}}$ $\mathbf{74.31}{\scriptstyle\pm{1.10}}$
$O(d)$-NSD $76.09{\scriptstyle\pm 1.79}$ $81.79{\scriptstyle\pm 1.63}$ $74.07{\scriptstyle\pm 1.58}$
Gen-BSNN $\textbf{{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}76.32}}{\scriptstyle\pm 1.65}$ $\textbf{{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}82.03}}{\scriptstyle\pm{0.94}}$ $\textbf{{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}72.29}}{\scriptstyle\pm{1.21}}$
Gen-NSD $72.97{\scriptstyle\pm 1.81}$ $77.25{\scriptstyle\pm 1.67}$ $70.16{\scriptstyle\pm 2.08}$
GCN $56.50{\scriptstyle\pm 0.81}$ $55.63{\scriptstyle\pm 1.13}$ $51.97{\scriptstyle\pm 0.49}$
@@
Table 2: Average test accuracy across all hyperparameter configurations and corresponding standard deviations.
Texas Wisconsin Cornell
Diag-BSNN $\textbf{{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}73.41}}{\scriptstyle\pm 2.84}$ $\textbf{{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}77.78}}{\scriptstyle\pm 3.32}$ $\textbf{{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}64.77}}{\scriptstyle\pm 5.51}$
Diag-NSD $71.48{\scriptstyle\pm 3.54}$ $74.85{\scriptstyle\pm 4.57}$ $62.18{\scriptstyle\pm 7.10}$
$\mathbf{SO(d)}$-BSNN $\textbf{{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}72.88}}{\scriptstyle\pm 3.63}$ $\textbf{{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}76.84}}{\scriptstyle\pm 4.68}$ $\textbf{{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}65.21}}{\scriptstyle\pm 6.84}$
$O(d)$-NSD $71.98{\scriptstyle\pm 3.28}$ $74.55{\scriptstyle\pm 5.03}$ $61.73{\scriptstyle\pm 6.70}$
Gen-BSNN $\textbf{{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}73.12}}{\scriptstyle\pm 2.41}$ $\textbf{{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}77.92}}{\scriptstyle\pm 3.23}$ $\textbf{{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}65.77}}{\scriptstyle\pm 4.87}$
Gen-NSD $71.34{\scriptstyle\pm 2.70}$ $74.64{\scriptstyle\pm 3.92}$ $62.20{\scriptstyle\pm 5.69}$
@@ 6 Conclusion and Future Directions
In this work, we have introduced Bayesian sheaf neural networks: novel sheaf neural networks in which the sheaf Laplacian is a random latent variable, and where the loss function used in training contains a Kullback-Leibler divergence regularization term. Not only does this allow one to quantify uncertainty of the model arising from the learned sheaf, but we show that for node classifications tasks with limited training data, our BSNN models outperform deterministic sheaf models. Additionally, our experiments suggest that the BSNN models are more robust to different choices of hyperparameters and weight initializations than the corresponding deterministic sheaf models.
@@ Acknowledgements
This work has been partially supported by the Army Research Laboratory Cooperative Agreement No. W911NF2120186, and STRONG ARL CA No W911NF-22-2-0139.
Appendix A Proof of Proposition 2.3
@@
In what follows, given a cellular sheaf $\mathcal{F}$ over a graph $G=(V,E)$ and initial conditions $X(0)$, let $X(t)$ denote the solution to Eq. (3). Note that the limit of $X(t)$ as $t\to\infty$ exists, and is equal to the orthogonal projection of $X(0)$ onto $\ker(\Delta_{\mathcal{F}})$ [34]. If $\mathcal{F}$ has $d$--dimensional stalks and $G$ has $n$ nodes, although $X(t)\in\mathbb{R}^{nd}$, it will be useful to equally regard $X(t)$ as an $n\times d$ matrix. With this in mind, let $X(t)_{u}$ denote the row of $X(t)$ corresponding to node $u$. We write $u\in V_{l}$ if node $u$ has label $l$. For a cellular sheaf $\mathcal{F}$ with $d$--dimensional stalks over a graph $G=(V,E)$ with labeling $\ell:V\to\mathcal{L}$ and initial conditions $X(0)$, we say that $\mathcal{F}$ linearly separates the classes of $(G,X(0),\ell)$ if for each label $l\in\mathcal{L}$, the sets $\{\lim_{t\to\infty}X(t)_{u}:u\in V_{l}\}\subseteq\mathbb{R}^{d}$ and $\{\lim_{t\to\infty}X(t)_{u}:u\notin V_{l}\}\subseteq\mathbb{R}^{d}$ are linearly separable. With this terminology, Definition 2.1 states that a collection of cellular sheaves $\mathcal{H}$ has linear separation power over a collection of labelled graphs $\mathcal{G}$, if for each graph $G=(V,E)\in\mathcal{G}$ with labelling $\ell:V\to\mathcal{L}$, there exists a cellular sheaf $\mathcal{F}\in\mathcal{H}$ over $G$ such that $\mathcal{F}$ linearly separates the classes of $(G,X(0),\ell)$ for almost all initial conditions $X(0)$, i.e. all but a measure zero subset of the initial conditions.
Proof of Proposition 2.3.
@@
Let $d\geq 1$ and let $d=4k+m$ for $40\leq m<$ 4. That is, $k=\lfloor\tfrac{d}{4}\rfloor$. Since proving the proposition for $C=7k$ also proves the cases $C<7k$, we may assume that $C=7k$. Let =$\mathcal{L}=\{1,2,\dots,C\}$ denote the set of class labels for the nodes of a graph $G=(V,E)\in\mathcal{G}$ and let $\ell:V\to\mathcal{L}$ denote the labelling of the nodes of $G$. For each $1\leq i\leq k$, let $J_{i}=\{7(i-1)+1,7(i-1)+2,\dots,7(i-1)+7\}$, let $\mathcal{L}_{i}=\{j:j\in J_{i}\}\cup\{0\}$, and let $\ell_{i}:V\to\mathcal{L}_{i}$ denote a relabeling of $G$ where for each $u\in V$, $f_{i}(u)=f(u)$ whenever $f(u)\in J_{i}$, and $f_{i}(u)=0$ otherwise. That is, the relabeling $\ell_{i}$ retains the seven labels in $J_{i}$, and replaces all other labels with a single distinct eighth label.
@@
By [6, Proposition 13], for each $1\leq i\leq k$ we may find a cellular sheaf $\mathcal{F}^{i}\in\mathcal{H}_{so}^{4}$ with $4$--dimensional stalks which linearly separates the eight classes of $(G,X(0),\ell_{i})$ for almost all initial conditions $X(0)\in\mathbb{R}^{4n}$. Let $\mathcal{F}$ denote the cellular sheaf with restriction maps $\mathcal{F}_{u\trianglelefteq e}=\mathcal{F}^{1}_{u\trianglelefteq e}\oplus\mathcal{F}^{2}_{u\trianglelefteq e}\oplus\cdots\oplus\mathcal{F}^{k}_{u\trianglelefteq e}\oplus I_{m}$. We claim that $\mathcal{F}$ linearly separates the classes of the original labelling $(G,X(0),\ell)$ for almost all initial conditions $X(0)\in\mathbb{R}^{n\times d}$.
@@
Given initial conditions $X(0)\in\mathbb{R}^{n\times d}$, for each $1\leq i\leq k$ let $X_{i}(0)\in\mathbb{R}^{n\times 4}$ be the matrix consisting of columns $34i-$ 3 through $4i$ of $X(0)$. Note that for almost all initial conditions $X(0)$, it will be true that for all $1\leq i\leq k$, $\mathcal{F}^{i}$ linearly separates the classes of $(G,X_{i}(0),\ell_{i})$. Next, for any class $l\in\mathcal{L}$, find $i$ such that $l\in J_{i}$. Since $\mathcal{F}^{i}$ linearly separates the classes of $(G,X_{i}(0),\ell_{i})$, we may find a hyperplane $A\subseteq\mathbb{R}^{4}$ which separates the class $l$ from all other classes in $\mathcal{L}_{i}$, and hence $\mathcal{L}$ as well. Then $\mathbb{R}^{4}\oplus\dots\oplus A\oplus\dots\oplus\mathbb{R}^{4}\oplus\mathbb{R}^{m}$ is a hyperplane in $\mathbb{R}^{d}$ separating class $l$ from all other classes. ∎
Appendix B Section 4 Proofs
@@
$\mathcal{H}^{d}_{\delta}(S)=\inf\Big{\{}\sum_{i=1}^{\infty}\alpha_{d}(\tfrac{1}{2}\operatorname{diam}(S))^{d}:\bigcup_{i=1}^{\infty}U_{i}\supseteq S,\operatorname{diam}(U_{i})<\delta\Big{\}}$
where $\alpha_{d}$ denotes the volume of a unit $d$--ball. The $d$--dimensional Hausdorff measure on $M$ is then given by $\mathcal{H}^{d}(S)=\lim_{\delta\to 0}H^{d}_{\delta}(S)$.
Let $\lambda^{n}$ denote the $n$--dimensional Lebesgue measure. The area formula [23] computes the Hausdorff measure of the image of a $1$-to-$1$ and continuously differentiable function $f$ in terms of the Jacobian of $f$. We state a version presented in [51].
Proposition B.1 (Area formula).
Let $f:\mathbb{R}^{n}\to\mathbb{R}^{m}$ be $1$-to-$1$ and continuously differentiable. For any Borel subset $A\subseteq\mathbb{R}^{n}$,

### v2

@@ 4.1 Results
Sheaf neural networks are expected to outperform traditional graph neural networks on datasets with low edge-homophily coefficients, that is, datasets where relatively few edges connect nodes belonging to the same class [69].
We evaluate our model using the NSD benchmark [10], which compiles standardized results across datasets with varying levels of homophily, from low (Texas) to high (Cora), enabling fair and diverse comparison. For more details on the datasets, see Appendix Appendix D: Experiment Details.
Baselines include standard GNNs designed for homophilic settings (GCN [48], GAT [62], GraphSAGE [38]), models tailored for heterophilic graphs (GGCN [63], Geom-GCN [57], H2GCN [69], GPRGNN [17], MixHop [1]), and architectures addressing oversmoothing (GCNII [16], PairNorm [68]). Most results are sourced from [63], with MixHop from [7] and [69].
Texas Wisconsin Film Cornell Citeseer Pubmed Cora
Diag-BSNN 85.95 $\pm$ 5.77 88.43 $\pm$ 2.39 37.22 $\pm$ 1.19 85.95 $\pm$ 6.71 76.91 $\pm$ 1.80 89.59 $\pm$ 0.42 87.34 $\pm$ 0.94
SO(d)-BSNN 85.95 $\pm$ 6.26 89.80 $\pm$ 3.49 37.20 $\pm$ 1.06 86.22 $\pm$ 6.10 76.80 $\pm$ 1.30 89.66 $\pm$ 0.42 86.68 $\pm$ 1.14
Gen-BSNN 88.11 $\pm$ 5.16 89.61 $\pm$ 4.21 37.31 $\pm$ 0.88 85.68 $\pm$ 6.40 77.17 $\pm$ 1.84 89.63 $\pm$ 0.42 87.53 $\pm$ 1.53
Diag-NSD 85.67 $\pm$ 6.95 88.63 $\pm$ 2.75 37.79 $\pm$ 1.01 86.49 $\pm$ 7.35 77.14 $\pm$ 1.85 89.42 $\pm$ 0.43 87.14 $\pm$ 1.06
O(d)-NSD 85.95 $\pm$ 5.51 89.41 $\pm$ 4.74 37.81 $\pm$ 1.15 84.86 $\pm$ 4.71 76.70 $\pm$ 1.57 89.49 $\pm$ 0.40 86.90 $\pm$ 1.13
Gen-NSD 82.97 $\pm$ 5.13 89.21 $\pm$ 3.84 37.80 $\pm$ 1.22 85.68 $\pm$ 6.51 76.32 $\pm$ 1.65 89.33 $\pm$ 0.35 87.30 $\pm$ 1.15
GGCN 84.86 $\pm$ 4.55 86.86 $\pm$ 3.29 37.54 $\pm$ 1.56 85.68 $\pm$ 6.63 77.14 $\pm$ 1.45 89.15 $\pm$ 0.37 87.95 $\pm$ 1.05
H2GCN 84.86 $\pm$ 7.23 87.65 $\pm$ 4.98 35.70 $\pm$ 1.00 82.70 $\pm$ 5.28 77.11 $\pm$ 1.57 89.49 $\pm$ 0.38 87.87 $\pm$ 1.20
GPRGNN 78.38 $\pm$ 4.36 82.94 $\pm$ 4.21 34.63 $\pm$ 1.22 80.27 $\pm$ 8.11 77.13 $\pm$ 1.67 87.54 $\pm$ 0.38 87.95 $\pm$ 1.18
MixHop 77.84 $\pm$ 7.73 75.88 $\pm$ 4.90 32.22 $\pm$ 2.34 73.51 $\pm$ 6.34 76.26 $\pm$ 1.33 85.31 $\pm$ 0.61 87.61 $\pm$ 0.85
GCNII 77.57 $\pm$ 3.83 80.39 $\pm$ 3.40 37.44 $\pm$ 1.30 77.86 $\pm$ 3.79 77.33 $\pm$ 1.48 90.15 $\pm$ 0.43 88.37 $\pm$ 1.25
Geom-GCN 66.76 $\pm$ 2.72 64.51 $\pm$ 3.66 31.59 $\pm$ 1.15 60.54 $\pm$ 3.67 78.02 $\pm$ 1.15 89.95 $\pm$ 0.47 85.35 $\pm$ 1.57
PairNorm 60.27 $\pm$ 4.34 48.43 $\pm$ 6.14 27.40 $\pm$ 1.24 58.92 $\pm$ 3.15 73.59 $\pm$ 1.47 87.53 $\pm$ 0.44 85.79 $\pm$ 1.01
GraphSAGE 82.43 $\pm$ 6.14 81.18 $\pm$ 5.56 34.23 $\pm$ 0.99 75.95 $\pm$ 5.01 76.04 $\pm$ 1.38 88.45 $\pm$ 0.50 86.90 $\pm$ 1.04
GCN 55.14 $\pm$ 5.16 51.76 $\pm$ 3.06 27.32 $\pm$ 1.10 60.54 $\pm$ 5.30 76.50 $\pm$ 1.36 88.42 $\pm$ 0.50 86.98 $\pm$ 1.27
GAT 52.16 $\pm$ 6.63 49.41 $\pm$ 4.09 27.44 $\pm$ 0.89 61.89 $\pm$ 5.05 76.55 $\pm$ 1.23 87.30 $\pm$ 1.10 86.33 $\pm$ 0.48
MLP 80.81 $\pm$ 4.75 85.29 $\pm$ 3.31 36.53 $\pm$ 0.70 81.89 $\pm$ 6.40 74.02 $\pm$ 1.90 87.16 $\pm$ 0.37 75.69 $\pm$ 2.00
Table 1: Test accuracies and standard deviations for node classification on datasets sorted in ascending order of homophily. The top three models are colored by rank. Our models are indicated as BSNN.
All models, including ours, are evaluated on 10 fixed splits comprising 48% training, 32% validation, and 20% testing data. We report mean accuracy and standard deviation across these splits. Model configurations are selected via random hyperparameter search, with test results reported at the epoch yielding the best validation performance (see Appendix Appendix D: Experiment Details Table 7 for the full hyperparameter list). The final model is chosen based on the highest validation score across all candidates.
The results in Table 1 demonstrate that the BSNN models consistently attain peak or near-peak test accuracies across a wide range of datasets. On heterophilic datasets such as Texas and Wisconsin, BSNN variants are competitive with state-of-the-art methods, with accuracies reaching 88.11% on Texas and 89.80% on Wisconsin. Similarly, on homophilic benchmarks such as Pubmed and Cora, BSNN variants match or slightly exceed advanced methods, achieving 89.66% and 87.53% accuracy, respectively. These results indicate that BSNN models are robust to varying degrees of homophily and generalize effectively across different graph structures.
Notably, on the challenging Film dataset with low homophily, sparse connectivity, and high label noise, BSNN models achieve test accuracy comparable to leading methods, indicating robustness to noisy and weakly informative graph structures.
Overall, these results demonstrate that the incorporation of Bayesian uncertainty modeling into graph neural networks provides significant benefits across a broad spectrum of datasets.
4.2 Uncertainty Quantification
Given that BSNN models learn predictive distributions on sheaves, we evaluate their reliability, and calibration using metrics such as entropy, epistemic variance, mutual information, and expected calibration error as outlined in Section 2.5. We conduct independent training trials using 30 random seeds with the best hyperparameter configuration, namely the one that achieved the highest validation accuracy. For each trial, a BSNN model samples a new sheaf instance and performs three forward passes. We use these predictions to compute node-wise uncertainty statistics.
For each test node, we calculate entropy to quantify total uncertainty, and estimate epistemic variance to capture model uncertainty. As shown in Figure 3, epistemic variance is more sensitive to model choice than entropy. This variability in epistemic uncertainty across models is natural and expected, as different models embody distinct prior beliefs and architectures, which interact with datasets in different ways. In contrast, entropy remains relatively stable across models for a given dataset, suggesting that it is likely dominated by aleatoric uncertainty.
Most nodes exhibit low epistemic variance (near 0) and low to moderate entropy, suggesting that models are generally confident and consistent in their predictions, with no clear correlation between epistemic variance and homophily level across models. The maximum observed epistemic variance (0.08) is on Cornell when Gen-BSNN is used, while the rest of the plots show a maximum epistemic variance of about 0.04. Node degree does not appear to significantly influence epistemic variance, as Citeseer and Film both exhibit low variance under Gen-BSNN despite having markedly different average degrees, where Citeseer being sparsely connected and Film more densely connected.
Among the three BSNN variants, Diag-BSNN exhibits lower epistemic variance than Gen-BSNN and $SO(d)$-BSNN in most datasets. This suggests that its simplified diagonal posterior often leads to more conservative uncertainty estimates, but its behavior still varies depending on the specific dataset characteristics.
In most datasets, models demonstrate confident predictions for the majority of nodes, as indicated by a high concentration of low-entropy values. Only a smaller subset of nodes shows elevated entropy and variability, which typically points to localized uncertainty arising from ambiguous features. In contrast, the Film dataset shows more consistent entropy levels across nodes, with overall low variance, suggesting that the observed uncertainty is likely due to noise rather than informative structure. This interpretation is supported by the low mutual information (MI) values, as will be shown in Figure 4.
Table 2 presents average uncertainty scores, computed by first averaging predictive entropy and epistemic variance within each dataset, and then averaging those values across all datasets. The observed averages suggest that all models are similarly confident with moderately low uncertainty overall.
Average Entropy Average Epistemic Variance
Diag-BSNN 0.474796 0.000381
Gen-BSNN 0.449049 0.001247
SO(d)-BSNN 0.525914 0.000738
Table 2: Average entropy and epistemic variance for BSNN models across all datasets.
Figure 3: Predictive entropy and epistemic variance for BSNN models, illustrating total and model uncertainty across nodes.
Figure 4: Mutual information for BSNN models as a measure of epistemic uncertainty.
Since predictive entropy is nearly invariant across models, we compute mutual information (MI) to better capture differences in uncertainty. The histograms that illustrate mutual information in Figure 4 are heavily skewed toward 1.0 across all datasets and all models. This suggests that the BSNN models are very confident about most of their predictions. As for Cornell, a small proportion of samples have a higher MI, meaning the model is uncertain about these specific samples, likely due to out-of-distribution features, or decision boundary proximity. This is not surprising, as these samples also exhibited the highest epistemic uncertainty, further indicating the model’s lack of confidence in its predictions for these cases.
We compute the Expected Calibration Error (ECE) to assess the alignment between predicted confidence and actual accuracy, and report the results in Table 3. On datasets with high homophily levels, the ECE values that are below 0.1 show that the predicted probabilities closely align with actual outcomes, and that the models are well-calibrated. In contrast, on heterophilic datasets such as Texas, Wisconsin, and Film, the ECE values are higher. This suggests that while the models may achieve high accuracy, they are overconfident. To address this, post-hoc calibration techniques such as temperature scaling could be applied to improve the reliability of confidence scores without retraining the model.
Texas Wisconsin Film Cornell Citeseer Pubmed Cora
Diag-BSNN 0.1483 0.0918 0.2490 0.1564 0.0391 0.0113 0.0473
SO(d)-BSNN 0.2061 0.1128 0.1168 0.2602 0.0401 0.0144 0.0296
Gen-BSNN 0.1998 0.1343 0.2652 0.2072 0.0440 0.0144 0.0449
Table 3: Expected Calibration Error (ECE) for different BSNN variants across datasets. Values below 0.1 are highlighted in blue to indicate well-calibrated models.
4.3 Robustness Under Limited Data
To evaluate the performance of our Bayesian sheaf models under limited training data, we use 10 fixed dataset splits, with each split allocating 32% of the nodes for training, 20% for validation, and 48% for testing. For each model configuration, we report test results from the epoch with the best validation score. The final model is selected as the configuration with the highest validation performance overall. As shown in Table 4, our model still achieves high accuracy in most datasets, even in this low-data regime.
Texas Wisconsin Film Cornell Citeseer Pubmed Cora
Diag-BSNN 80.11 ± 5.24 83.25 ± 5.16 36.83 ± 0.74 74.60 ± 3.19 73.71 ± 1.04 88.74 ± 0.26 84.52 ± 1.25
SO(d)-BSNN 79.31 ± 7.27 84.00 ± 3.96 35.47 ± 0.78 74.37 ± 4.99 73.95 ± 2.59 88.92 ± 0.29 83.98 ± 1.06
Gen-BSNN 79.66 ± 5.34 83.33 ± 4.62 36.83 ± 0.56 74.25 ± 5.22 73.98 ± 1.18 88.72 ± 0.36 84.50 ± 1.12
Table 4: Test accuracies and standard deviations for node classification under limited training data conditions, using the hyperparameter configuration yielding the best validation accuracy.
Additionally, to examine the sensitivity of the Bayesian and deterministic sheaf models to the hyperparameter selections, we perform a grid search over selected hyperparameters, as detailed in Appendix Appendix D: Experiment Details, Table 8. We report the mean test accuracy across all hyperparameter configurations in Table 5.
@@ 4.3 Robustness Under Limited Data
The advantage of each BSNN model over the corresponding NSD model when considering the entire range of hyperparameters is definitive. A Wilcoxon signed-rank test for a $p$-value threshold of $p=0.05$ indicates a statistically significant improvement in performance for the Diag-BSNN and Gen-BSNN models over their NSD counterparts across datasets.
Thus, we observe that the BSNN models maintain good performance over a much wider range of hyperparameters when compared with the NSD model with the same type of restriction maps. We also include results for a graph convolutional network (GCN) as a baseline comparison.
Texas Wisconsin Film Cornell Citeseer Pubmed Cora
Diag-BSNN 73.48 ± 2.65 77.56 ± 3.60 35.36 ± 1.02 65.17 ± 5.71 68.53 ± 3.97 87.76 ± 0.59 76.14 ± 5.13
SO(d)-BSNN 72.75 ± 3.51 76.96 ± 4.59 34.51 ± 1.26 64.90 ± 6.69 68.34 ± 4.91 87.88 ± 0.63 76.09 ± 5.40
Gen-BSNN 73.73 ± 2.67 78.05 ± 3.25 35.43 ± 0.85 66.35 ± 4.53 69.08 ± 3.56 87.81 ± 0.53 76.63 ± 4.59
Diag-NSD 71.22 ± 3.31 74.91 ± 4.69 34.99 ± 1.31 62.43 ± 7.11 66.59 ± 5.76 88.01 ± 0.58 75.26 ± 7.85
O(d)-NSD 71.86 ± 3.16 74.63 ± 5.07 34.63 ± 1.26 61.87 ± 6.38 65.14 ± 6.37 88.03 ± 0.63 73.04 ± 9.65
Gen-NSD 71.44 ± 2.47 74.24 ± 3.91 35.11 ± 1.00 61.93 ± 5.79 64.17 ± 6.85 87.87 ± 0.48 73.48 ± 10.5
GCN 55.11 ± 1.42 50.23 ± 2.81 27.22 ± 1.77 52.27 ± 0.39 70.19 ± 2.22 86.01 ± 0.99 84.21 ± 1.24
@@
Table 5: Average test accuracy and corresponding standard deviations across all hyperparameter configurations, evaluated under limited training data conditions.
@@ 5 Conclusion and Future Directions
In this work, we have introduced Bayesian sheaf neural networks: novel sheaf neural networks in which the sheaf Laplacian is a random latent variable, and where the loss function used in training contains a Kullback-Leibler divergence regularization term. Not only does this allow one to quantify uncertainty of the model arising from the learned sheaf, but we show that our BSNN models achieve results that are comparable to, and in some cases exceed, state-of-the-art node classification methods. Furthermore, our experiments suggest that the BSNN models are more robust to different choices of hyperparameters and weight initializations, particularly under limited training data, compared to their deterministic counterparts.
@@ Acknowledgments
This work has been partially supported by the Army Research Laboratory Cooperative Agreement No W911NF2120186, and STRONG ARL CA No W911NF-22-2-0139.
Appendix A: Proof of Proposition 1
@@
In what follows, given a cellular sheaf $\mathcal{F}$ over a graph $G=(V,E)$ and initial conditions $X(0)$, let $X(t)$ denote the solution to Eq.(1). Note that the limit of $X(t)$ as $t\to\infty$ exists, and is equal to the orthogonal projection of $X(0)$ onto $\ker(\Delta_{\mathcal{F}})$ [41]. If $\mathcal{F}$ has $d$-dimensional stalks and $G$ has $n$ nodes, although $X(t)\in\mathbb{R}^{nd}$, it will be useful to equally regard $X(t)$ as an $n\times d$ matrix.
Let $X(t)_{u}$ denote the row of $X(t)$ corresponding to node $u$. We write $u\in V_{l}$ if node $u$ has label $l$. For a cellular sheaf $\mathcal{F}$ with $d$-dimensional stalks over a graph $G=(V,E)$ with labeling $\ell:V\to\mathcal{L}$ and initial conditions $X(0)$, we say that $\mathcal{F}$ linearly separates the classes of $(G,X(0),\ell)$ if for each label $l\in\mathcal{L}$, the sets $\{\lim_{t\to\infty}X(t)_{u}:u\in V_{l}\}\subseteq\mathbb{R}^{d}$ and $\{\lim_{t\to\infty}X(t)_{u}:u\notin V_{l}\}\subseteq\mathbb{R}^{d}$ are linearly separable.
With this terminology, Definition 6 states that a collection of cellular sheaves $\mathcal{H}$ has linear separation power over a collection of labeled graphs $\mathcal{G}$, if for each graph $G=(V,E)\in\mathcal{G}$ with labeling $\ell:V\to\mathcal{L}$, there exists a cellular sheaf $\mathcal{F}\in\mathcal{H}$ over $G$ such that $\mathcal{F}$ linearly separates the classes of $(G,X(0),\ell)$ for almost all initial conditions $X(0)$, i.e. all but a measure zero subset of the initial conditions.
Proof of Proposition 1.
@@
Let $d\geq 1$ and let $d=4k+m$ for $0\leq m<4$. That is, $k=\lfloor\tfrac{d}{4}\rfloor$. Since proving the proposition for $C=7k$ also proves the cases $C<7k$, we may assume that $C=7k$. Let $\mathcal{L}=\{1,2,\dots,C\}$ denote the set of class labels for the nodes of a graph $G=(V,E)\in\mathcal{G}$ and let $\ell:V\to\mathcal{L}$ denote the labeling of the nodes of $G$.
For each $1\leq i\leq k$, let $J_{i}=\{7(i-1)+1,7(i-1)+2,\dots,7(i-1)+7\}$, let $\mathcal{L}_{i}=\{j:j\in J_{i}\}\cup\{0\}$, and let $\ell_{i}:V\to\mathcal{L}_{i}$ denote a relabeling of $G$ where for each $u\in V$, $f_{i}(u)=f(u)$ whenever $f(u)\in J_{i}$, and $f_{i}(u)=0$ otherwise. That is, the relabeling $\ell_{i}$ retains the seven labels in $J_{i}$, and replaces all other labels with a single distinct eighth label.
@@
According to Proposition 13 in [10], for each $1\leq i\leq k$ we may find a cellular sheaf $\mathcal{F}^{i}\in\mathcal{H}_{so}^{4}$ with 44-dimensional stalks which linearly separates the eight classes of $(G,X(0),\ell_{i})$ for almost all initial conditions $X(0)\in\mathbb{R}^{4n}$. Let $\mathcal{F}$ denote the cellular sheaf with restriction maps $\mathcal{F}_{u\trianglelefteq e}=\mathcal{F}^{1}_{u\trianglelefteq e}\oplus\mathcal{F}^{2}_{u\trianglelefteq e}\oplus\cdots\oplus\mathcal{F}^{k}_{u\trianglelefteq e}\oplus I_{m}$. We claim that $\mathcal{F}$ linearly separates the classes of the original labeling $(G,X(0),\ell)$ for almost all initial conditions $X(0)\in\mathbb{R}^{n\times d}$.
@@
Given initial conditions $X(0)\in\mathbb{R}^{n\times d}$, for each $1\leq i\leq k$ let $X_{i}(0)\in\mathbb{R}^{n\times 4}$ be the matrix consisting of columns 4i34i-3 through $4i$ of $X(0)$. Note that for almost all initial conditions $X(0)$, it will be true that for all $1\leq i\leq k$, $\mathcal{F}^{i}$ linearly separates the classes of $(G,X_{i}(0),\ell_{i})$. Next, for any class $l\in\mathcal{L}$, find $i$ such that $l\in J_{i}$. Since $\mathcal{F}^{i}$ linearly separates the classes of $(G,X_{i}(0),\ell_{i})$, we may find a hyperplane $A\subseteq\mathbb{R}^{4}$ which separates the class $l$ from all other classes in $\mathcal{L}_{i}$, and hence $\mathcal{L}$ as well. Then $\mathbb{R}^{4}\oplus\dots\oplus A\oplus\dots\oplus\mathbb{R}^{4}\oplus\mathbb{R}^{m}$ is a hyperplane in $\mathbb{R}^{d}$ separating class $l$ from all other classes. ∎
Appendix B: Section 3.3 Proofs
@@
$\mathcal{H}^{d}_{\delta}(S)=\inf\Big{\{}\sum_{i=1}^{\infty}\alpha_{d}(\tfrac{1}{2}\operatorname{diam}(S))^{d}:\bigcup_{i=1}^{\infty}U_{i}\supseteq S,\operatorname{diam}(U_{i})<\delta\Big{\}},$
where $\alpha_{d}$ denotes the volume of a unit $d$-ball. The $d$-dimensional Hausdorff measure on $M$ is then given by $\mathcal{H}^{d}(S)=\lim_{\delta\to 0}H^{d}_{\delta}(S)$.
Let $\lambda^{n}$ denote the $n$-dimensional Lebesgue measure. The area formula in [29] computes the Hausdorff measure of the image of a 11-to-11 and continuously differentiable function $f$ in terms of the Jacobian of $f$. We state a version presented in [60].
Proposition 4 (Area formula).
Let $f:\mathbb{R}^{n}\to\mathbb{R}^{m}$ be 11-to-11 and continuously differentiable. For any Borel subset $A\subseteq\mathbb{R}^{n}$,